The length of a rectangular garden is 9 feet longer than its width. If the garden's perimeter is 202 feet, what is the area of the garden in square feet?
Answer by lwsshak3(11628)

The length of a rectangular garden is 9 feet longer than its width. If the garden's perimeter is 202 feet, what is the area of the garden in square feet?
***
let x=width
x+9=length
Area=length*width=x(x+9)
perimeter=2x+2(x+9)=202
x+x+9=101
2x+9=101
2x=92
x=46(width)
x+9=55(length)
area=x(x+9)=46*55=2530 sq ft